SOLVE THE FOLLOWING EQUATIONS
15+20=2x+2x
6x+2x=20-4
5x-17=30x+3
3(x+2-3x)=20+2(x-2)
10(x+5-3)=5x-3
4x-10+30=4x+7x-8
5x-10+2(3+x-2x)=8x
3x-2=4(-2+3x)+20

Answer by jcmtnez(53)

Solving
15+20=2x+2x, 35=4x, x=35/4
6x+2x=20-4, 8x=16, x=2
5x-17=30x+3, -20=25x, x=-4/5
3(x+2-3x)=20+2(x-2), 3x+6-9x=20+2x-4, -10=8x, -5/4=x
10(x+5-3)=5x-3, 10x+50-30=5x-3, 5x=-23, x=-23/5
4x-10+30=4x+7x-8, 28=7x, 4=x
5x-10+2(3+x-2x)=8x, 5x-10+6+2x-4x=8x, -4=5x, -4/5=x
3x-2=4(-2+3x)+20, 3x-2=-8+12x+20, -14=9x, x=-14/9